# Coordinate space representation of a one dimensional $p$-wave pseudopotential

### v1

@@ Abstract
We propose a discrete-space representation of the $p$-wave pseudopotential. The proposed representation is validated by applying it to the analytically solvable case of two fermions in a harmonic trap and successfully recovering the exact energy spectrum and eigenfunctions. Furthermore, we use the square-well and modified Pöschl–Teller potentials as finite-range representations of the $p$-wave interaction and study their convergence to the contact interaction when the range tends to zero. Finally, we perform natural orbital analysis and compute the eigenvalues of the one-body density matrix for different particle numbers, examining their dependence on the one-dimensional scattering length and identifying distinct physical regimes.
@@ I Introduction
Quantum fluctuations are enhanced as dimensionality is reduced, being stronger in 2D and even more pronounced in 1D. One-dimensional systems, in particular, are often more tractable for numerical treatment and may even admit exact solutions [11, 12]. Significant advances in trapping and cooling techniques [13, 14] have enabled the confinement of cold atomic gases in a one-dimensional ($1D$) regime, achieved using waveguides with tight transverse confinement [15, 16]. The quasi-$1D$ regime is realized when the chemical potential and the thermal energy are smaller than the energy level spacing of the transverse oscillator [17]. Under these conditions, the longitudinal motion prevails over other directions, and the system can effectively be considered a one-dimensional harmonic oscillator.
@@
In ultracold quantum gases, interactions are generally dominated by ss‑wave scattering [18]. However, in systems of identical fermions with spin fully polarized, the Pauli exclusion principle forces the spatial wave function to be antisymmetric, which completely forbids ss‑wave contact interactions. Consequently, the leading interaction channel becomes pp‑wave scattering. This has been observed and explored both experimentally and theoretically, for instance, in measurements of a $p$-wave Feshbach resonance in ultracold $^{40}K$ gases [19] and in studies of $p$-wave resonant threshold behavior [20]. Subsequent work has investigated $p$-wave collisions in low-dimensional geometries [21], and more recent theoretical and quantum Monte Carlo studies have characterized the roles of the $p$-wave scattering volume and effective range in spin-polarized Fermi systems [22]. Motivated by this body of work, we analyze properties of a system of $N$ spin-aligned fermions in a $1D$ harmonic trap, where the spin state is frozen in the fully polarized configuration $\uparrow_{1}\uparrow_{2}\ldots\uparrow_{N}$, and pp‑wave scattering is the only allowed short‑range contact interaction.
@@
The structure of this work is as follows. In Sec. II, we introduce the Hamiltonian of the system and show two limits for which analytical solutions are known. In Sec. III, we derive a representation of the $p$-wave interaction in discrete space, which reproduces $p$-wave scattering in the zero-range limit. In Sec. IV, we present two possible representations of the $p$-wave interaction in continuous space using two different potentials: a square well and the modified Pöschl–Teller potential. This framework allows us to model $p$-wave interaction as a function of the interaction range. In Sec. V, we present the relative energy spectrum of a two-fermion system, using the different representations of the interaction derived in the previous sections. Moreover, we study the one-body density matrix for different scattering lengths and numbers of fermions. Finally, in Sec. VI, we summarize and present the main conclusions of our work.
@@ II Model
Consider $N$ spin-aligned fermions confined in a $1D$ harmonic trap and interacting via $p$-wave scattering. The Hamiltonian of the system is [23]
$\hat{\mathcal{H}}=\sum_{i=1}^{N}-\frac{\hbar^{2}}{2m}\partial_{x_{i}}^{2}+\frac{1}{2}m\omega^{2}x_{i}^{2}-\sum_{j>k}^{N}g_{\text{F}}\delta^{\prime}(x_{kj})\hat{\partial}_{kj},$ (1)
@@
The one-dimensional scattering length, $a_{1D}$, is defined as the position of the node, closest to $x=0$, of the analytic continuation from large distances of the two-body scattering solution,
$\psi(x)\propto x-a_{1D},\quad x\to\infty\;.$ (3)
The zero-range interaction potential, $g_{\text{F}}\delta^{\prime}(x_{i}-x_{j})\hat{\partial}_{ij}$, is characterized by the coupling constant related to the scattering length as $g_{\text{F}}=2a_{1D}\hbar^{2}/m$. Thus, for a zero-range potential, the scattering length is related to the Bethe-Peierls condition for the logarithmic derivative of the wave function at the origin,
@@
The use of a contact interaction potential for $p$-wave fermions is therefore problematic, as it would require the wave function to be discontinuous as it must satisfy the boundary condition (4), while also remaining antisymmetric to obey Fermi-Dirac statistics. Indeed, contact interaction acts only when two fermions occupy the same position, while the Pauli exclusion principle forbids such situation. As a result, $p$-wave interactions cannot be properly described by zero-range interaction potentials and must instead be introduced as the appropriate limit of finite-range interactions.
Analytic solutions for a general number of particles are known in two limits, the non-interacting case and the infinitely interacting limit. In the latter, the system is solvable by mapping the original Hamiltonian to that of an ideal Bose gas [24]. The ground state energy of infinitely interacting $p$-wave fermions is equal to that of non-interacting bosons. In this limit, the system is known as the fermionic Tonks-Girardeau (FTG) gas. This mapping is analogous to the well-known correspondence between an impenetrable Bose gas and an ideal Fermi gas [25, 26], where the ground state energy of non-interacting fermions matches that of infinitely repulsive bosons. Outside these limits, the $p$-wave problem must be solved numerically.
@@ III Discrete representation
In this section, we propose an explicit discrete-space representation of the $p$-wave interaction. In this framework, the $s$-wave interaction, $g_{\text{B}}\delta(x_{i}-x_{j})$, is typically represented by a potential of depth $g_{\text{B}}/\Delta x$ at mesh points where $x_{i}=x_{j}$, with $\Delta x$ denoting the spatial discretization step. We extend this concept by introducing a potential that accurately captures the physics of the $p$-wave interaction.
To derive the discrete representation, we take advantage of the existence of an exact solution to a $p$-wave problem in the continuous representation and invert the stationary Schrödinger equation to express the interaction potential $V$ in terms of the second derivative of the eigenfunction $\Psi(x_{1},\ldots,x_{N})$,
$V(x_{1},\ldots,x_{N})=$ $E+\frac{\hbar^{2}}{2m}\frac{\sum_{i=1}^{N}\partial_{x_{i}}^{2}\Psi(x_{1},\ldots,x_{N})}{\Psi(x_{1},\ldots,x_{N})}$ (5)
$-U(x_{1},\ldots,x_{N}),$
@@
We consider the system of two spin-aligned fermions in free space with $p$-wave interaction, for which the solutions are known analytically. The Hamiltonian of the system is
@@
$\psi_{\text{rel}}(x)=\sqrt{\frac{\hbar^{2}}{2\mu g}}\text{sgn}(x)e^{-|x|\hbar^{2}/(\mu g_{\text{F}})},$ (7)
with energy $E_{\text{rel}}=-\hbar^{6}/(2\mu^{3}g_{\text{F}}^{2})$ and $\mu=m/2$ is the reduced mass. This solution is used to determine $V$ from Eq. (5).
To perform the calculation numerically in position space, we discretize it. The choice of mesh directly determines the shape of the potential. At $x=0$, the sign function is equal to $\text{sgn}(x=0)=0$, and Eq. (5) gives $V(x=0)\to\infty$, introducing a divergent Hamiltonian term that complicates numerical diagonalization. To avoid this issue, the mesh is defined as $\{-L,-L+\Delta x,\ldots,-\Delta x,\Delta x,\ldots,L-\Delta x,L\}$, explicitly excluding $x=0$. Within this grid, the discrete representation is non-zero only at the points $x=\pm\Delta x$, with height equal to
$V(\pm\Delta x)=$ $\frac{\hbar^{2}}{2\mu}\frac{\psi_{\text{rel}}(0)-2\psi_{\text{rel}}(-\Delta x)+\psi_{\text{rel}}(-2\Delta x)}{\Delta x^{2}\psi_{\text{rel}}(-\Delta x)}$ (8)
$+E_{\text{rel}}=\frac{\hbar^{2}}{\mu\Delta x^{2}}\left(\frac{e^{-\gamma}}{2}-1-\frac{\gamma^{2}}{2}\right),$
where $\psi_{\text{rel}}(0)=0$, and $\gamma\equiv\Delta x\hbar^{2}/(\mu g_{\text{F}})$. Figure 1 shows the resulting potential as a function of the interparticle separation. Note that $x=0$ is absent in the mesh.
In summary, we have introduced an explicit discrete space representation of the $p$-wave interaction valid in the zero-range limit ($R\to 0$)
$V(x_{1},\ldots)=\sum_{i<j}^{N}\frac{\hbar^{2}}{\mu\Delta x^{2}}\left(\frac{e^{-\gamma}}{2}-1-\frac{\gamma^{2}}{2}\right)\delta_{x_{i},x_{j}\pm\Delta x},$ (9)
where $\delta_{x_{i},x_{j}\pm\Delta x}$ is the Kronecker delta. This interaction is non-zero only at $x_{i}=x_{j}\pm\Delta x$. If additional points are placed between the points $x_{i}=x_{j}\pm\Delta x$, the representation would no longer be valid. Therefore, the interaction must be defined such that it is non-zero only at $x_{i}=x_{j}\pm\Delta x$, with no intermediate points. Moreover, this formulation is derived assuming that the discrete second derivative is computed as
@@ IV Continuous representation
In this section, we study two continuous-space representations of the $p$-wave interaction using two different potentials: a square well and the modified Pöschl–Teller potential [28]. We focus on the case of two spin-polarized fermions in free space and determine how the potential depths relate to the scattering length and the interaction range, $R$.
IV.1 Square well potential
@@
$V_{SW}(x)=\left\{\begin{aligned} -&V_{0}&\quad&\text{if}&|x|<R_{SW}\\ &0&\quad&\text{if}&|x|\geq R_{SW}\end{aligned}\right$.\,, (11)
@@
$\left\{\begin{aligned} &\partial^{2}_{x}\psi(x)+\kappa^{2}\psi(x)=0&\quad&\text{if}&|x|<R_{SW}\\ &\partial^{2}_{x}\psi(x)=0&\quad&\text{if}&|x|\geq R_{SW}\end{aligned}\right$.\,, (12)
with $\kappa=\sqrt{2\mu V_{0}/\hbar^{2}}$ being the characteristic wavenumber related to the depth of the interaction potential. Due to fermionic antisymmetry, the solution must have odd parity. The zero-energy scattering solution to the Schrödinger equation is
$\left\{\begin{aligned} \psi(x)&=N_{2}(x+a_{1D})&\quad&\text{if}&&x\leq-R_{SW}\\ \psi(x)&=N_{1}\sin(\kappa x)&\quad&\text{if}&|&x|<R_{SW}\\ \psi(x)&=N_{2}(x-a_{1D})&\quad&\text{if}&&x\geq R_{SW}\end{aligned}\right$.\,, (13)
@@
We assume that $\kappa R_{SW}<4.493$, ensuring that $\psi(x)$ has only a single node at $x=0$ inside the well. The effective range for the square well potential is
@@ IV.2 Modified Pöschl–Teller potential
Figure 1: Discrete and continuous representations of the $p$-wave interaction as a function of the relative coordinate between fermions, $x$. The discrete representation is shown as red crosses, with no specific point defined at $x=0$ (its value is irrelevant for the antisymmetric $p$-wave solutions). The square well potential, $V_{SW}$, is depicted by a dashed blue line while the modified Pöschl–Teller potential, $V_{PT}$, is shown by a dash-dotted green line.
@@
$V_{PT}(x)=-\frac{\hbar^{2}}{2\mu R_{PT}^{2}}\frac{\lambda(\lambda-1)}{\cosh^{2}(x/R_{PT})},$ (16)
where $\lambda>1$ is a dimensionless parameter defining the interaction strength. While formally potential (16) has an infinite range, the distance $R_{PT}$ effectively defines its physical interaction range. Figure 1 illustrates the shape of the modified Pöschl–Teller potential. For positive energies and at large positive distance, the asymptotic behavior of the odd-parity eigenfunctions is given by [30]
$\psi(x)\to$ $\frac{\sqrt{\pi}}{2}\Bigg{\{}\frac{\Gamma\left(-ikR_{PT}\right)e^{ik\ln(2)R_{PT}}e^{-ik|x|}}{\Gamma\left(\frac{\lambda+1}{2}-\frac{ikR_{PT}}{2}\right)\Gamma\left(1-\frac{\lambda}{2}-\frac{ikR_{PT}}{2}\right)}$ (17)
$+\text{c.c.}\Bigg{\}},$
where $k^{2}=2\mu E/\hbar^{2}$. Expanding Eq. (17) for $k\to 0$, a linear series decomposition allows one to locate the node of the wave function and thereby obtain the $1D$ scattering length

### v2

# Coordinate space representation of a one-dimensional odd-parity pseudopotential

@@ Abstract
We propose a discrete-space representation of a one-dimensional zero-range odd-parity pseudopotential. The proposed representation is validated by applying it to the analytically solvable case of two fermions in a harmonic trap and successfully recovering the exact energy spectrum and eigenfunctions. Furthermore, we use the square-well and modified Pöschl–Teller potentials as finite-range representations of the odd-parity interaction and study their convergence to the contact interaction when the range tends to zero. Finally, we perform natural orbital analysis and compute the eigenvalues of the one-body density matrix for different particle numbers, examining their dependence on the one-dimensional scattering length and identifying distinct physical regimes.
@@ I Introduction
Quantum fluctuations are enhanced as dimensionality is reduced, being stronger in $2D$ and even more pronounced in $1D$. One-dimensional systems, in particular, are often easier to handle numerically and may even admit exact solutions [12, 32]. Significant advances in trapping and cooling techniques [28, 25] have enabled the confinement of cold atomic gases in a one-dimensional regime, achieved using waveguides with tight transverse confinement [6, 17]. The quasi-$1D$ regime is realized when the chemical potential and the thermal energy are smaller than the energy level spacing of the transverse oscillator [26]. Under these conditions, the longitudinal motion prevails over other directions, and the system can effectively be considered a one-dimensional harmonic oscillator.
@@
In ultracold quantum gases, interactions are generally dominated by ss‑wave scattering [33]. However, in systems of identical fermions with spin fully polarized, the Pauli exclusion principle forces the spatial wave function to be antisymmetric, which completely forbids ss‑wave contact interactions. Consequently, the leading interaction channel becomes pp‑wave scattering. This has been observed and explored both experimentally and theoretically, for instance, in measurements of a $p$-wave Feshbach resonance in ultracold $^{40}K$ gases [29] and in studies of $p$-wave resonant threshold behavior [9]. Subsequent work has investigated $p$-wave collisions in low-dimensional geometries [Günter2005], and more recent theoretical and quantum Monte Carlo studies have characterized the roles of the $p$-wave scattering volume and effective range in spin-polarized Fermi systems [2]. Motivated by this body of work, we analyze properties of a system of $N$ spin-aligned fermions in a $1D$ harmonic trap, where the spin state is frozen in the fully polarized configuration $\uparrow_{1}\uparrow_{2}\ldots\uparrow_{N}$, and pp‑wave scattering is the only allowed short‑range contact interaction.
In one-dimensional systems, the notion of a $p$-wave interaction must be reformulated, since the absence of angular degrees of freedom precludes the usual partial-wave expansion familiar from three dimensions. Instead, what is conventionally referred to as a $p$-wave interaction in $1D$ is more precisely captured by an odd-parity pseudopotential, which enforces the appropriate symmetry of the relative wave function under particle exchange. This odd-parity contact interaction reproduces the low-energy scattering properties associated with the $3D$ $p$-wave channel when the system is strongly transversely confined, giving rise to effective $1D$ scattering parameters such as the scattering length and effective range.
@@
The structure of this work is as follows. In Sec. II, we introduce the Hamiltonian of the system and show two limits for which analytical solutions are known. In Sec. III, we derive a representation of the odd-parity interaction in discrete space, which reproduces $p$-wave scattering in the zero-range limit. In Sec. IV, we present two possible representations of the interaction in continuous space using two different potentials: a square-well and the modified Pöschl–Teller potential. This framework allows us to model the odd-parity interaction as a function of the interaction range. In Sec. V, we present the relative energy spectrum of a two-fermion system, using the different representations of the interaction derived in the previous sections. Moreover, we study the one-body density matrix for different scattering lengths and numbers of fermions. Finally, in Sec. VI, we summarize and present the main conclusions of our work.
@@ II Model
Consider $N$ spin-aligned fermions confined in a $1D$ harmonic trap and interacting via an odd-parity zero-range pseudopotential. The Hamiltonian of the system is [15]
$\hat{\mathcal{H}}=\sum_{i=1}^{N}-\frac{\hbar^{2}}{2m}\partial_{x_{i}}^{2}+\frac{1}{2}m\omega^{2}x_{i}^{2}-g_{\text{F}}\sum_{j>k}^{N}\delta^{\prime}(x_{kj})\hat{\partial}_{kj},$ (1)
@@
The one-dimensional scattering length, $a_{1D}$, is defined as the position of the node, closest to $x_{kj}=0$, obtained by extrapolating the asymptotic linear behavior of the two-body scattering wave function
$\psi(x_{kj})\propto x_{kj}-a_{1D},\quad x_{kj}\to\infty.$ (3)
The zero-range interaction potential, $g_{\text{F}}\delta^{\prime}(x_{kj})\hat{\partial}_{kj}$, is characterized by the coupling constant related to the scattering length as $g_{\text{F}}=2a_{1D}\hbar^{2}/m$. Thus, for a zero-range potential, the scattering length is related to the Bethe-Peierls condition for the logarithmic derivative of the wave function at the origin,
@@
The use of a simple delta-function contact potential for $p$-wave fermions is problematic, as it cannot reproduce the correct boundary condition Eq. (4) while preserving antisymmetry under particle exchange. Since a $\delta$-potential acts only when two fermions occupy the same position, it has no effect on antisymmetric wave functions that vanish at contact. Consequently, odd-parity interactions cannot be directly modeled using such simple contact potentials, including those typically employed in exact-diagonalization methods. A proper description requires either derivative (slope-dependent) zero-range potentials or the appropriate finite-range limit, which might be more involved to implement numerically.
Analytic solutions for a general number of particles are known in two limits, the non-interacting case and the infinitely interacting limit. In the latter, the system is solvable by mapping the original Hamiltonian to that of an ideal Bose gas [13]. The ground state energy of infinitely interacting fermions is equal to that of non-interacting bosons. In this limit, the system is known as the fermionic Tonks-Girardeau (FTG) gas. This mapping is analogous to the well-known correspondence between an impenetrable Bose gas and an ideal Fermi gas [14, 34], where the ground state energy of non-interacting fermions matches that of infinitely repulsive bosons. Outside these limits, no analytic solution is known for $N>2$, and the problem must therefore be addressed numerically.
@@ III Discrete representation
In this section, we propose an explicit discrete-space representation of the odd-parity interaction in the zero-range limit. In this framework, the $s$-wave interaction, $g_{\text{B}}\delta(x_{kj})$, is typically represented by a potential of depth $g_{\text{B}}/\Delta x$ at mesh points where $x_{i}=x_{j}$, with $\Delta x$ denoting the spatial discretization step. We extend this concept by introducing a potential that accurately captures the physics of the odd-parity interaction.
To derive the discrete representation, we take advantage of the existence of an exact solution to a fermionic problem in the continuous representation and invert the stationary Schrödinger equation to express the interaction potential $V$ in terms of the second derivative of the eigenfunction $\Psi(x_{1},\ldots,x_{N})$,
$aligned V(x_{1},\ldots,x_{N})=&E+\frac{\hbar^{2}}{2m}\frac{\sum_{i=1}^{N}\partial_{x_{i}}^{2}\Psi(x_{1},\ldots,x_{N})}{\Psi(x_{1},\ldots,x_{N})}\\ &-U(x_{1},\ldots,x_{N}),$ (5)
@@
We consider the system of two spin-aligned fermions in free space with odd-parity zero-range interaction, for which the solutions are known analytically. The Hamiltonian of the system is
@@
$\psi_{\text{rel}}(x)=\sqrt{\frac{\hbar^{2}}{mg_{\text{F}}}}\text{sgn}(x)e^{-2|x|\hbar^{2}/(mg_{\text{F}})},$ (7)
with energy $E_{\text{rel}}=-4\hbar^{6}/(m^{3}g_{\text{F}}^{2})$. This solution is used to determine $V$ from Eq. (5).
To perform the calculation numerically in position space, we discretize it. The choice of mesh directly determines the shape of the potential. At $x=0$, the sign function is equal to $\text{sgn}(x=0)=0$, and Eq. (5) gives $V(x=0)\to\infty$, introducing a divergent Hamiltonian term that prevents numerical diagonalization. To avoid this issue, the mesh is defined as $\{-L,-L+\Delta x,\ldots,-\Delta x,\Delta x,\ldots,L-\Delta x,L\}$, explicitly excluding $x=0$. Within this grid, the discrete representation is non-zero only at the points $x=\pm\Delta x$, with height equal to
[…]
where $\psi_{\text{rel}}(0)\equiv 0$, and $\gamma\equiv 2\Delta x\hbar^{2}/(mg_{\text{F}})$. Figure 1 shows the resulting potential as a function of the interparticle separation. Note that $x=0$ is absent in the mesh.
In summary, we have introduced an explicit discrete space representation of the odd-parity interaction valid in the zero-range limit ($R\to 0$)
$V(x_{1},\ldots)=\sum_{i<j}^{N}\frac{\hbar^{2}}{m\Delta x^{2}}\left(e^{-\gamma}-2-\gamma^{2}\right)\delta_{x_{i},x_{j}\pm\Delta x},$ (9)
where $\delta_{x_{i},x_{j}\pm\Delta x}$ is the Kronecker delta. This interaction is non-zero only at $x_{i}=x_{j}\pm\Delta x$. Moreover, this formulation is derived assuming that the discrete second derivative is computed as
@@
The Hamiltonian with potential Eq. (8) is invariant under the exchange of any two particles, leading to $N!$ degenerate energy levels. To obtain the fermionic states, we introduce a symmetry-breaking term in the Hamiltonian that assigns an energy penalty to eigenstates lacking the required antisymmetry, thereby restricting the low-energy manifold to purely fermionic wave functions. This additional term appears in the off-diagonal elements of the Hamiltonian as
$split V_{\mathrm{odd}}\left[\psi(x_{1},\ldots)\right]&=K\sum_{P\in\mathcal{S}_{N}}(-1)^{\epsilon(P)}\psi(x_{P_{1}},\ldots),$ (11)
where $\mathcal{S}_{N}$ denotes the Symmetric group, and $(-1)^{\epsilon(P)}$ reflects the parity of the permutation $P$. If $P$ can be written as an even number of transpositions, it evaluates to +1+1; if an odd number of transpositions is required, it evaluates to 1-1. This term contributes an energy of $N!K$ to the fermionic states. By choosing $|K|\gg 1$ and $K<0$, all the fermionic solutions are shifted to the low-energy sector. As a result, the physically relevant eigenvalues can be obtained simply by computing the smallest eigenvalues of the full Hamiltonian.
Although Eq. (9) is originally derived for two fermions in free space with an odd-parity interaction and a positive scattering length, it yields accurate results for harmonically trapped systems for both positive and negative scattering lengths.
@@ IV Continuous representation
In this section, we study two continuous-space representations of the odd-parity interaction using two different potentials: a square-well and the modified Pöschl–Teller potential [10]. We focus on the case of two spin-polarized fermions in free space and determine how the potential depths relate to the scattering length and the interaction range, $R$.
IV.1 Square-well potential
@@
[…]
@@
[…]
with $\kappa=\sqrt{mV_{0}/\hbar^{2}}$ being the characteristic wavenumber related to the depth of the interaction potential. Due to fermionic antisymmetry, the solution must have odd parity. The zero-energy scattering solution to the Schrödinger equation is
[…]
@@
We assume that $\kappa R_{SW}<4.493$, limiting the maximum value of the potential depth, and ensuring that $\psi(x)$ has only a single node at $x=0$ inside the well. The effective range for the square-well potential is
@@ IV.2 Modified Pöschl–Teller potential
Figure 1: Discrete and continuous representations of the odd-parity interaction as a function of the relative coordinate between fermions, $x$. The discrete representation is shown as red crosses, with no specific point defined at $x=0$. The square-well potential, $V_{SW}$, is depicted by a dashed blue line while the modified Pöschl–Teller potential, $V_{PT}$, is shown by a dash-dotted green line.
@@
$V_{PT}(x)=-\frac{\hbar^{2}}{mR_{PT}^{2}}\frac{\lambda(\lambda-1)}{\cosh^{2}(x/R_{PT})},$ (17)
where $\lambda>1$ is a dimensionless parameter defining the interaction strength. While formally potential Eq. (17) has an infinite range, the distance $R_{PT}$ effectively defines its physical interaction range. Figure 1 illustrates the shape of the modified Pöschl–Teller potential. For positive energies and at large positive distance, the asymptotic behavior of the odd-parity eigenfunctions is given by [11]
[…]
where $k^{2}=mE/\hbar^{2}$. Expanding Eq. (18) for $k\to 0$, a linear series decomposition allows one to locate the node of the wave function and thereby obtain the $1D$ scattering length